# JEL classification: C31, J24

### v1

@@ 3.2 Identification and Estimation
We now discuss how the asymptotic variance of $\boldsymbol{\hat{\psi}}$ can be consistently estimated. One simple approach is to use the robust estimator à la White (1980) that allows for the variance of the component of $\boldsymbol{\eta}_{s}$ and $\boldsymbol{\varepsilon}_{s}$ to vary across schools. Let $\mathbf{\tilde{X}}_{s}=\mathbf{J}_{s}[\mathbf{X}_{s},\leavevmode\nobreak\ \mathbf{G}_{s}\mathbf{X}_{s}]$, $\mathbf{R}_{s}=[\mathbf{J}_{s}\mathbf{G}_{s}\mathbf{y}_{s},\leavevmode\nobreak\ \mathbf{\tilde{X}}_{s}]$, $\mathbf{Z}_{s}=[\mathbf{J}_{s}\mathbf{G}_{s}^{2}\mathbf{X}_{s},\leavevmode\nobreak\ \mathbf{\tilde{X}}_{s}]$, $\mathbf{R}^{\prime}\mathbf{Z}=\sum_{s=1}^{S}\mathbf{R}^{\prime}_{s}\mathbf{Z}_{s}$, $\mathbf{Z}^{\prime}\mathbf{Z}=\sum_{s=1}^{S}\mathbf{Z}^{\prime}_{s}\mathbf{Z}_{s}$. Let also $\hat{\boldsymbol{v}}_{s}:=\mathbf{J}_{s}\mathbf{y}_{s}-\mathbf{R}_{s}\boldsymbol{\hat{\psi}}$, that is, $\hat{\boldsymbol{v}}_{s}$ is the residual vector from Equation (7). We denote by $\operatorname{diag}$ the bloc diagonal matrix operator. The asymptotic variance of $\boldsymbol{\hat{\psi}}$ can be estimated by $\mathbf{\hat{B}}_{n}^{-1}\mathbf{\hat{D}}_{n}\mathbf{\hat{B}}_{n}^{-1}/n$, where $\mathbf{\hat{B}}_{n}=(\mathbf{R}^{\prime}\mathbf{Z})(\mathbf{Z}^{\prime}\mathbf{Z})^{-1}(\mathbf{R}^{\prime}\mathbf{Z})^{\prime}/n$ and $\mathbf{\hat{D}}_{n}=(\mathbf{R}^{\prime}\mathbf{Z})(\mathbf{Z}^{\prime}\mathbf{Z})^{-1}\operatorname{diag}\{\mathbf{Z}_{1}^{\prime}\hat{\boldsymbol{v}}_{1}\hat{\boldsymbol{v}}_{1}^{\prime}\mathbf{Z}_{1},\dots,\mathbf{Z}_{S}^{\prime}\hat{\boldsymbol{v}}_{S}\hat{\boldsymbol{v}}_{S}^{\prime}\mathbf{Z}_{S}\}(\mathbf{Z}^{\prime}\mathbf{Z})^{-1}(\mathbf{R}^{\prime}\mathbf{Z})^{\prime}/n$.
@@
(i) $\mathbf{G}_{s}$is uniformly bounded in column sum; (ii) $\lambda\neq 0$; (iii) $(\eta_{s,i},\leavevmode\nobreak\ \varepsilon_{s,i})$is independently distributed across $i$ such that $\mathbb{E}(\eta_{s,i}^{2}|\mathbf{G}_{s},\mathbf{X}_{s})=\sigma_{\eta}^{2}>0$, $\mathbb{E}((\delta^{2}\varepsilon_{s,i})^{2}|\mathbf{G}_{s},\mathbf{X}_{s})=\sigma_{\epsilon}^{2}>0$, $\mathbb{E}(\delta^{2}\varepsilon_{s,i}\eta_{s,i}|\mathbf{G}_{s},\mathbf{X}_{s})=\rho\sigma_{\eta}\sigma_{\epsilon}$, and $\lvert\rho\rvert\leq 1$.
If both $S$ and $n_{s}$ tend to infinity, Condition (i) rules out the cases where the sum of certain columns of $\mathbf{G}_{s}$ is unbounded. This condition is also considered by (Lee, 2004) in the case of the standard model. Condition (iii) imposes constant variances for the error terms $\eta_{s,i}$ and $\varepsilon_{s,i}$ but accounts for their potential correlation. This is important as $\eta_{s,i}$ and $\varepsilon_{s,i}$ characterize the same student $i$. The parameter $\rho$ measures the correlation between $\eta_{s,i}$ and $\varepsilon_{s,i}$. We impose no restrictions on the joint distribution of $(\eta_{s,i},\leavevmode\nobreak\ \varepsilon_{s,i})$. The restriction $\lambda\neq 0$ of Condition (ii) is necessary for identifying $\sigma_{\eta}$, $\sigma_{\epsilon}$, and $\rho$. If $\lambda=0$, the disturbance of Equation (7) would be $\mathbf{J}_{s}\boldsymbol{\eta}_{s}+\delta^{2}\mathbf{J}_{s}\boldsymbol{\varepsilon}_{s}$, and one cannot disentangle $\sigma_{\eta}$, $\sigma_{\epsilon}$, and $\rho$.
Under Assumption 3.3, we can estimate $(\sigma_{\eta}^{2},\leavevmode\nobreak\ \sigma_{\epsilon}^{2},\leavevmode\nobreak\ \rho)$ and construct a consistent estimator for $\operatorname{\mathbb{V}}(\boldsymbol{\hat{\psi}})$. We employ a quasi-maximum likelihood (QML) approach, where the dependent variable is the vector of the residual vector $\hat{\boldsymbol{v}}_{s}=\mathbf{J}_{s}\mathbf{y}_{s}-\mathbf{R}_{s}\boldsymbol{\hat{\psi}}$. The likelihood of $\hat{\boldsymbol{v}}_{s}$ is based on a multivariate normal distribution with mean and covariance matrix equal to those of the true error term $\boldsymbol{v}_{s}$, where $\lambda$ is replaced by its estimator $\hat{\lambda}$. Importantly, we do not require $\boldsymbol{v}_{s}$ to be actually normally distributed. Let $(\hat{\sigma}_{\eta}^{2},\leavevmode\nobreak\ \hat{\sigma}_{\epsilon}^{2},\leavevmode\nobreak\ \hat{\rho})$ be the QML estimator of $(\sigma_{\varepsilon}^{2},\leavevmode\nobreak\ \sigma_{\epsilon}^{2},\leavevmode\nobreak\ \rho)$. We establish the following result.
@@
Under Proposition 3.1 and Assumptions A.2A.4 stated in Appendix A.4, $(\sigma_{\epsilon}^{2},\leavevmode\nobreak\ \sigma_{\epsilon}^{2},\leavevmode\nobreak\ \rho)$ is globally identified, and $(\hat{\sigma}_{\eta}^{2},\leavevmode\nobreak\ \hat{\sigma}_{\epsilon}^{2},\leavevmode\nobreak\ \hat{\rho})$ is a consistent estimator.
We can now consistently estimate the asymptotic variance of $\boldsymbol{\hat{\psi}}$ using the estimator $(\hat{\sigma}_{\eta}^{2},\leavevmode\nobreak\ \hat{\sigma}_{\epsilon}^{2},\leavevmode\nobreak\ \hat{\rho})$. We denote by $\mathbf{B}=\operatorname{plim}(\mathbf{R}^{\prime}\mathbf{Z})(\mathbf{Z}^{\prime}\mathbf{Z})^{-1}(\mathbf{R}^{\prime}\mathbf{Z})^{\prime}/n$, where $\operatorname{plim}$ stands for the limit in probability as $n$ grows to infinity. We also define $\boldsymbol{\check{\Omega}}=\sum_{s=1}^{S}\mathbf{Z}_{s}^{\prime}\big{(}\sigma_{\epsilon}^{2}\mathbf{J}_{s}+\sigma_{\eta}^{2}\mathbf{J}_{s}^{\prime}\mathbf{W}_{s}\mathbf{W}_{s}^{\prime}\mathbf{J}_{s}+\rho\sigma_{\epsilon}\sigma_{\eta}\mathbf{J}_{s}(\mathbf{W}_{s}+\mathbf{W}_{s}^{\prime})\mathbf{J}_{s})\big{)}\mathbf{Z}_{s}$. We make the assumption that $\operatorname{plim}(\mathbf{R}^{\prime}\mathbf{Z})(\mathbf{Z}^{\prime}\mathbf{Z})^{-1}\boldsymbol{\check{\Omega}}(\mathbf{Z}^{\prime}\mathbf{Z})^{-1}(\mathbf{R}^{\prime}\mathbf{Z})^{\prime}/n$ exists and is denoted by $\mathbf{D}$. As a result $\lim_{n\to\infty}n\operatorname{\mathbb{V}}(\boldsymbol{\hat{\psi}})=\mathbf{B}^{-1}\mathbf{D}\mathbf{B}^{-1}$. We can obtain a consistent estimator of the variance by replacing $\sigma_{\epsilon}^{2}$, $\sigma_{\eta}^{2}$, and $\rho$ with their estimator, and $\mathbf{B}$ and $\mathbf{D}$ with their empirical counterparts.
@@ 3.3 Simulation Study
[…]
@@
We consider four estimation approaches for each DGP. The most flexible estimation approach follows Equation (6), which can be expressed as follows:
@@
$y_{s,i}=\kappa^{NI}_{s}\ell^{NI}_{s,i}+\kappa^{I}_{s}(1-\ell^{NI}_{s,i})+\mathbf{g}_{s,i}\mathbf{y}_{s}+\mathbf{x}^{\prime}_{s,i}\boldsymbol{\tilde{\beta}}+\mathbf{g}_{s,i}\mathbf{X}_{s}\boldsymbol{\tilde{\gamma}}+\tilde{v}_{s,i},$ (9)
@@
where $\ell^{NI}_{s,i}$ is a dummy variable indicating whether the student is not isolated, $\kappa^{I}_{s}=\delta^{2}c_{s}+\alpha_{s}$, $\kappa^{NI}_{s}=\delta^{2}c_{s}+(1-\lambda)\alpha_{s}$, and $\tilde{v}_{s,i}=(\boldsymbol{\omega}_{s,i}-\lambda\mathbf{g}_{s,i})\boldsymbol{\eta}_{s}+\delta^{2}\varepsilon_{s,i}$. This specification, referred to as Model 4, is our proposed model. Additionally, we estimate several variants of this specification that impose constraints on $c_{s}$ and $\alpha_{s}$. These constrained specifications align with those commonly estimated in the literature when approximating students’ efforts by their GPA:
  • Model 1: Specification (9) without school fixed effects which assumes that $\kappa^{I}_{s}=\kappa^{NI}_{s}=\kappa$, where $\kappa$ is a constant ($c_{s}$ does not vary across schools and $\alpha_{s}=0$).
@@
  • Model 2: Specification (9) with school fixed effects, which assumes that $\kappa^{I}_{s}=\kappa^{NI}_{s}=\kappa_{s}$, where $\kappa_{s}$ varies across schools ($c_{s}$ varies across schools, whereas $\alpha_{s}=0$).
  • Model 3: Specification (9) with school fixed effects and a dummy variable for isolated students, which assumes that $\kappa^{NI}_{s}$ varies across schools and $\kappa^{I}_{s}=\kappa$, where $\kappa$ is a constant. This means that $\alpha_{s}=\bar{\alpha}$ may not be zero but does not vary across schools, whereas $c_{s}$ varies across schools.
Table 1 presents a summary of our Monte Carlo results across 1,000 replications. Model 1 yields biased estimates for all GPAs, which is expected due to the absence of fixed effects in this model. Although the bias of the peer effect estimate is not substantial for DGP A, estimates for other parameters, particularly $\tilde{\beta}_{2}$ and $\tilde{\gamma}_{2}$, exhibit notable biases. This is because DGP A features preference shocks $c_{s}$, generated from $\mathbf{x}_{2s}$, which Model 1 fails to account for.
@@
Model 2 is the standard model commonly employed in the literature, especially when GPA approximates effort. This model demonstrates strong performance with DGP A, despite the presence of 21% isolated students. Interestingly, this result suggests that the dummy variable for isolated students might not be relevant, even when the network includes students without friends. As this dummy variable accounts for GPA shocks without influencing effort, its absence in DGP A (because $\alpha_{s}=0$) does not compromise the validity of estimates. However, Model 2 leads to biased estimates for DGPs B and C, where $\alpha_{s}\neq 0$. Importantly, these biases carry substantial policy implications. For instance, while the true value of the social multiplier effect, measured by $1/(1-\lambda)$ for non-isolated students, stands at 3.33, Model 2 underestimates it at 1.93 for DGP B and 1.73 for DGP C.
Model 3 incorporates a dummy variable for isolated students but assumes a constant coefficient associated with this variable across all schools. Adding this dummy variable means that $\alpha_{s}$ is a constant that may not be zero. This resolves the bias observed for DGP B but still results in biased estimates for DGP D. This is because DGP D assumes variability in $\alpha_{s}$ across schools.
@@
Model 4, which is the most flexible specification, performs well across all DGPs. The simulation results also suggest that we can consistently estimate the covariance structure parameters $\sigma^{2}_{\epsilon}$, $\sigma^{2}_{\eta}$, and $\rho$, as stated in Proposition 3.2. Yet, it is worth noting that the standard deviations of the estimates for DGPs A and B are higher in Model 4 compared to Model 3. This is because both models are suitable for these DGPs, whereas Model 4 is overparameterized. One can address this issue by comparing Model 3 to Model 4 using a Hausman specification test (Hausman, 1978). If the null hypothesis that the parameter $\boldsymbol{\psi}$ is the same for Models 3 and 4 is not rejected, then Model 3 would be preferred over Model 4.
@@
Model 1 Model 2 Model 3 Model 4
Mean Sd Mean Sd Mean Sd Mean Sd
@@
$\lambda=0.7$ $0.727$ $0.011$ $0.700$ $0.012$ $0.700$ $0.012$ $0.701$ $0.018$
$\tilde{\beta}_{1}=1$ $0.917$ $0.061$ $1.001$ $0.047$ $1.001$ $0.047$ $1.001$ $0.047$
$\tilde{\beta}_{2}=1.5$ $0.633$ $0.110$ $1.500$ $0.083$ $1.500$ $0.083$ $1.499$ $0.083$
$\tilde{\gamma}_{1}=5$ $5.018$ $0.099$ $4.999$ $0.091$ $5.000$ $0.096$ $5.003$ $0.098$
$\tilde{\gamma}_{2}=-3$ $-3.542$ $0.109$ $-3.004$ $0.090$ $-3.002$ $0.124$ $-2.997$ $0.172$
$\sigma^{2}_{\epsilon}=8$ $9.354$ $5.266$ $9.122$ $5.216$
$\sigma^{2}_{\eta}=15$ $15.752$ $6.105$ $15.329$ $6.024$
$\rho=0.4$ $0.448$ $0.513$ $0.459$ $0.511$
@@
$\lambda=0.7$ $0.569$ $0.035$ $0.483$ $0.033$ $0.700$ $0.012$ $0.701$ $0.018$
$\tilde{\beta}_{1}=1$ $1.655$ $0.151$ $0.979$ $0.103$ $1.001$ $0.047$ $1.001$ $0.047$
$\tilde{\beta}_{2}=1.5$ $2.413$ $0.286$ $1.424$ $0.174$ $1.500$ $0.083$ $1.499$ $0.083$
$\tilde{\gamma}_{1}=5$ $3.862$ $0.312$ $3.719$ $0.313$ $5.000$ $0.096$ $5.003$ $0.098$
$\tilde{\gamma}_{2}=-3$ $-8.343$ $0.413$ $-7.673$ $0.524$ $-3.002$ $0.124$ $-2.997$ $0.172$
$\sigma^{2}_{\epsilon}=8$ $9.354$ $5.266$ $9.122$ $5.216$
$\sigma^{2}_{\eta}=15$ $15.752$ $6.105$ $15.329$ $6.024$
$\rho=0.4$ $0.448$ $0.513$ $0.459$ $0.511$

### v2

@@ 3.2 Identification and Estimation
We employ a proof by contradiction. The nonidentification issue arises when $\mathbb{E}(\mathbf{J}_{s}\mathbf{G}_{s}\mathbf{y}_{s}|\mathbf{G}_{s},\mathbf{X}_{s})$ is perfectly collinear with $\mathbf{J}_{s}\mathbf{X}_{s}$ and $\mathbf{J}_{s}\mathbf{G}_{s}\mathbf{X}_{s}$. For a non-isolated student, this suggests that there exist vector of parameters, $\dot{\boldsymbol{\beta}}$ and $\dot{\boldsymbol{\gamma}}$, such that $\mathbb{E}(\mathbf{g}_{s,i}\mathbf{y}_{s}-y^{NI}_{s}|\mathbf{G}_{s},\mathbf{X}_{s})=(\mathbf{x}_{s,i}-\mathbf{\hat{x}}_{s})^{\prime}\dot{\boldsymbol{\beta}}+(\mathbf{g}_{s,i}\mathbf{X}_{s}-\bar{\mathbf{x}}^{NI}_{s})^{\prime}\dot{\boldsymbol{\gamma}}$, where $y^{NI}_{s}$, $\mathbf{x}^{NI}_{s}$, and $\bar{\mathbf{x}}^{NI}_{s}$ are respectively the averages of $\mathbf{g}_{s,i}\mathbf{y}_{s}$, $\mathbf{x}_{s,i}$, and $\mathbf{g}_{s,i}\mathbf{X}_{s}$ in $\mathcal{V}^{NI}_{s}$. The variables $\mathbf{g}_{s,i}\mathbf{y}_{s}$, $\mathbf{x}_{s,i}$, and $\mathbf{g}_{s,i}\mathbf{X}_{s}$ are taken in deviation with respect to their average in $\mathcal{V}^{NI}_{s}$ because of the matrix $\mathbf{J}_{s}$ that multiplies the terms of Equation (7). If we take the previous equation in difference between students $i_{1}$ and $i_{3}$, we obtain
We now show that Condition (i) of Assumption 3.2 is not compatible with Equation (8). As $i_{3}$ is $i_{1}$’s only friend and $i_{4}$ is $i_{3}$’s only friend, we have $\mathbf{g}_{s,i_{1}}\mathbf{y}_{s}=y_{s,i_{3}}$, $\mathbf{g}_{s,i_{3}}\mathbf{y}_{s}=y_{s,i_{4}}$, $\mathbf{g}_{s,i_{1}}\mathbf{X}_{s}=\mathbf{x}_{s,i_{3}}$, and $\mathbf{g}_{s,i_{3}}\mathbf{X}_{s}=\mathbf{x}_{s,i_{4}}$. Thus, Equation (8) implies $y_{i_{3}}^{e}-y_{i_{4}}^{e}=(\mathbf{x}_{s,i_{1}}-\mathbf{x}_{s,i_{3}})^{\prime}\dot{\boldsymbol{\beta}}+(\mathbf{x}_{s,i_{3}}-\mathbf{x}_{s,i_{4}})^{\prime}\dot{\boldsymbol{\gamma}}$, where $y_{i}^{e}=\mathbb{E}(y_{s,i}|\mathbf{G}_{s},\mathbf{X}_{s})$. Since $\mathbf{x}_{s,i_{2}}$ does not appear in the previous equation, it follows that, an increase in $\mathbf{x}_{s,i_{2}}$ (keeping $\mathbf{x}_{s,i_{1}}$, $\mathbf{x}_{s,i_{3}}$, and $\mathbf{x}_{s,i_{4}}$ fixed) has no impact on $y_{i_{3}}^{e}-y_{i_{4}}^{e}$. Put differently, this increase has either no impact on $y_{i_{3}}^{e}$ and $y_{i_{4}}^{e}$ or the same impact on both $y_{i_{3}}^{e}$ and $y_{i_{4}}^{e}$. The first implication is not possible because $\mathbf{x}_{s,i_{2}}$ is the vector of contextual variables for $i_{4}$ and Condition (i) of Assumption 3.2 implies that GPA is influenced by at least one contextual variable. Moreover, $\mathbf{x}_{s,i_{2}}$ cannot influence $y_{i_{3}}^{e}$ and $y_{i_{4}}^{e}$ in the same way because $i_{2}$ is a direct friend of $i_{4}$ and not directly linked to $i_{3}$. Consequently, $\mathbb{E}(\mathbf{J}_{s}\mathbf{G}_{s}\mathbf{y}_{s}|\mathbf{G}_{s},\mathbf{X}_{s})$ cannot be a linear combination of $\mathbf{J}_{s}\mathbf{X}_{s}$ and $\mathbf{J}_{s}\mathbf{G}_{s}\mathbf{X}_{s}$.
As in the case of the standard model, the regressor $\mathbf{J}_{s}\mathbf{G}_{s}\mathbf{y}_{s}$ is endogenous in Equation (7). However, it can be instrumented by $\mathbf{J}_{s}\mathbf{G}_{s}^{2}\mathbf{X}_{s}$ because $\mathbf{J}_{s}\mathbf{G}_{s}^{2}\mathbf{X}_{s}$ is an excluded regressor with is correlated to $\mathbf{J}_{s}\mathbf{G}_{s}\mathbf{y}_{s}$(BramoulleDjebbariFortin2009). Let $\boldsymbol{\hat{\psi}}$ be the IV estimator of $\boldsymbol{\psi}$ using the instrument vector $\mathbf{J}_{s}\mathbf{G}_{s}^{2}\mathbf{X}_{s}$. [^12] To avoid a weak instrument issue, the pool of instruments can also be expanded to $\mathbf{J}_{s}[\mathbf{G}_{s}^{2}\mathbf{X}_{s},\penalty 10000\ \dots,\penalty 10000\ \mathbf{G}_{s}^{p}\mathbf{X}_{s}]$ for some integer $p>2$. We have the following result.
Proposition 3.1.
Under Assumptions 2.13.2 and LABEL:ass:app:ident:gmm (stated in Appendix LABEL:append:ident:reflection), $\boldsymbol{\psi}$ is globally identified, $\boldsymbol{\hat{\psi}}$ is a consistent estimator, and $)\displaystyle\sqrt{n}(\boldsymbol{\hat{\psi}}-\boldsymbol{\psi}_{0})\overset{d}{\to}\mathcal{N}\big(0,\lim_{n\to\infty}n\operatorname{\mathbb{V}}(\boldsymbol{\hat{\psi}}$$\big)$, where $\boldsymbol{\psi}_{0}$ is the true value of $\boldsymbol{\psi}$.
The proof of Proposition 3.1 is presented in Appendix LABEL:append:ident:reflection. The consistency and asymptotic normality of $\boldsymbol{\hat{\psi}}$ are directly derived from kelejian1998generalized.
@@
We now discuss how the asymptotic variance of $\boldsymbol{\hat{\psi}}$ can be consistently estimated. One simple approach is to use the robust estimator à la white1980heteroskedasticity that allows for the variance of the component of $\boldsymbol{\eta}_{s}$ and $\boldsymbol{\varepsilon}_{s}$ to vary across schools. Let $\mathbf{\tilde{X}}_{s}=\mathbf{J}_{s}[\mathbf{X}_{s},\penalty 10000\ \mathbf{G}_{s}\mathbf{X}_{s}]$, $\mathbf{R}_{s}=[\mathbf{J}_{s}\mathbf{G}_{s}\mathbf{y}_{s},\penalty 10000\ \mathbf{\tilde{X}}_{s}]$, $\mathbf{Z}_{s}=[\mathbf{J}_{s}\mathbf{G}_{s}^{2}\mathbf{X}_{s},\penalty 10000\ \mathbf{\tilde{X}}_{s}]$, $\mathbf{R}^{\prime}\mathbf{Z}=\sum_{s=1}^{S}\mathbf{R}^{\prime}_{s}\mathbf{Z}_{s}$, $\mathbf{Z}^{\prime}\mathbf{Z}=\sum_{s=1}^{S}\mathbf{Z}^{\prime}_{s}\mathbf{Z}_{s}$. Let also $\hat{\boldsymbol{v}}_{s}:=\mathbf{J}_{s}\mathbf{y}_{s}-\mathbf{R}_{s}\boldsymbol{\hat{\psi}}$, that is, $\hat{\boldsymbol{v}}_{s}$ is the residual vector from Equation (7). We denote by $\operatorname{diag}$ the bloc diagonal matrix operator. The asymptotic variance of $\boldsymbol{\hat{\psi}}$ can be estimated by $\mathbf{\hat{B}}_{n}^{-1}\mathbf{\hat{D}}_{n}\mathbf{\hat{B}}_{n}^{-1}/n$, where $\mathbf{\hat{B}}_{n}=(\mathbf{R}^{\prime}\mathbf{Z})(\mathbf{Z}^{\prime}\mathbf{Z})^{-1}(\mathbf{R}^{\prime}\mathbf{Z})^{\prime}/n$ and $\mathbf{\hat{D}}_{n}=(\mathbf{R}^{\prime}\mathbf{Z})(\mathbf{Z}^{\prime}\mathbf{Z})^{-1}\operatorname{diag}\{\mathbf{Z}_{1}^{\prime}\hat{\boldsymbol{v}}_{1}\hat{\boldsymbol{v}}_{1}^{\prime}\mathbf{Z}_{1},\dots,\mathbf{Z}_{S}^{\prime}\hat{\boldsymbol{v}}_{S}\hat{\boldsymbol{v}}_{S}^{\prime}\mathbf{Z}_{S}\}(\mathbf{Z}^{\prime}\mathbf{Z})^{-1}(\mathbf{R}^{\prime}\mathbf{Z})^{\prime}/n$.
@@
(i) $\mathbf{G}_{s}$is uniformly bounded in column sum; (ii) $\lambda\neq 0$; (iii) $(\eta_{s,i},\penalty 10000\ \varepsilon_{s,i})$is independently distributed across $i$ such that $\mathbb{E}(\eta_{s,i}^{2}|\mathbf{G}_{s},\mathbf{X}_{s})=\sigma_{\eta}^{2}>0$, $\mathbb{E}(\varepsilon_{s,i}^{2}|\mathbf{G}_{s},\mathbf{X}_{s})=\sigma_{\epsilon}^{2}>0$, $\mathbb{E}(\varepsilon_{s,i}\eta_{s,i}|\mathbf{G}_{s},\mathbf{X}_{s})=\rho\sigma_{\eta}\sigma_{\epsilon}$, and $\lvert\rho\rvert\leq 1$.
If both $S$ and $n_{s}$ tend to infinity, Condition (i) rules out the cases where the sum of certain columns of $\mathbf{G}_{s}$ is unbounded. This condition is also considered by (lee2004asymptotic) in the case of the standard model. Condition (iii) imposes constant variances for the error terms $\eta_{s,i}$ and $\varepsilon_{s,i}$ but accounts for their potential correlation. This is important as $\eta_{s,i}$ and $\varepsilon_{s,i}$ characterize the same student $i$. The parameter $\rho$ measures the correlation between $\eta_{s,i}$ and $\varepsilon_{s,i}$. We impose no restrictions on the joint distribution of $(\eta_{s,i},\penalty 10000\ \varepsilon_{s,i})$. The restriction $\lambda\neq 0$ of Condition (ii) is necessary for identifying $\sigma_{\eta}$, $\sigma_{\epsilon}$, and $\rho$. If $\lambda=0$, the disturbance of Equation (7) would be $\mathbf{J}_{s}\boldsymbol{\eta}_{s}+\mathbf{J}_{s}\boldsymbol{\varepsilon}_{s}$, and one cannot disentangle $\sigma_{\eta}$, $\sigma_{\epsilon}$, and $\rho$.
[…]
@@
[…]
@@ 3.3 Simulation Study
We consider two exogenous variables in the matrix $\mathbf{X}_{s}$: $\mathbf{x}_{1s}\sim N(E_{1s},\penalty 10000\ 16)$ and $\mathbf{x}_{2s}\sim\text{Poisson}(E_{2s})$, where $E_{1s}$, $E_{2s}$ are fixed for each school and drawn from a Uniform distribution over [0, $0,\penalty 10000\ 1$0]. The distribution of these two control variables is specific to each school.
We investigate three data-generating processes (DGPs) denoted A, B, and C. For all DGPs, we assign the following values to the parameters: $\lambda=0.7$, $\boldsymbol{\beta}=(1,\penalty 10000\ 1.5)^{\prime}$, $\boldsymbol{\gamma}=(5,\penalty 10000\ -3)^{\prime}$, $\sigma^{2}_{\epsilon}=8$, $\sigma^{2}_{\eta}=15$, and $\rho=0.4$. We assume that $(\eta_{s,i},\penalty 10000\ \varepsilon_{s,i})$ follows a bivariate normal distribution. We consider the same school preference shock for all DGPs, defined as $c_{s}=-1.5q_{90}(\mathbf{x}_{2s})$, where $q_{a}(.)$ represents the $a$-th percentile. Generating $c_{s}$ in this way ensures that $c_{s}$ is a fixed effect because it is dependent on the control variables. To illustrate how the nature of the shock $\alpha_{s}$ can affect the results, we vary the definition of $\alpha_{s}$ for each DGP. For DGP A, we set $\alpha_{s}=0$ for all $s$. For DGP C, we define $\alpha_{s}=10q_{90}(\mathbf{x}_{1s})$. For DGP B, we set $\alpha_{s}$ to the average of the values of $\alpha_{1},\penalty 10000\ \dots,\penalty 10000\ \alpha_{S}$ obtained for DGP C. For both DGPs A and B, $\alpha_{s}$ does not vary across schools. The only difference is that $\alpha_{s}=0$ for DGP A, whereas $\alpha_{s}\neq 0$ for DGP B.
@@
We consider three estimation approaches for each DGP. The most flexible estimation approach follows Equation (6), which can be expressed as follows:
@@
$y_{s,i}=\kappa^{NI}_{s}\ell^{NI}_{s,i}+\kappa^{I}_{s}(1-\ell^{NI}_{s,i})+\lambda\mathbf{g}_{s,i}\mathbf{y}_{s}+\mathbf{x}^{\prime}_{s,i}\boldsymbol{\beta}+\mathbf{g}_{s,i}\mathbf{X}_{s}\boldsymbol{\gamma}+\tilde{v}_{s,i},$ (9)
@@
where $\ell^{NI}_{s,i}$ is a dummy variable indicating whether the student is not isolated, $\kappa^{I}_{s}=c_{s}+\alpha_{s}$, $\kappa^{NI}_{s}=c_{s}+(1-\lambda)\alpha_{s}$, and $\tilde{v}_{s,i}=\eta_{s,i}-\lambda\mathbf{g}_{s,i}\boldsymbol{\eta}_{s}+\varepsilon_{s,i}$. This specification, referred to as Model 3, is our proposed model. Additionally, we estimate variants of this specification that impose constraints on $c_{s}$ and $\alpha_{s}$. These constrained specifications align with those commonly estimated in the literature when approximating students’ efforts by their GPA:
@@
  • Model 1: Specification (9) with school fixed effects, which assumes that $\kappa^{I}_{s}=\kappa^{NI}_{s}=\kappa_{s}$, where $\kappa_{s}$ varies across schools.
  • Model 2: Specification (9) with one school fixed effect $\kappa_{s}$ shared by all students in school $s$ plus a single global indicator for being isolated with coefficient $\phi$. Equivalently, $\kappa_{s}^{NI}=\kappa_{s}$ and $\kappa_{s}^{I}=\kappa_{s}+\phi$.
Table 1 presents a summary of our Monte Carlo results across 1,000 replications.
@@
Model 1 is the standard model commonly employed in the literature, especially when GPA approximates effort. This model demonstrates strong performance with DGP A, despite the presence of isolated students. Interestingly, this result suggests that the dummy variable for isolated students might not be relevant, even when the network includes students without friends. As this dummy variable accounts for GPA shocks without influencing effort, its absence in DGP A (because $\alpha_{s}=0$) does not compromise the validity of estimates. However, Model 1 leads to biased estimates for DGPs B and C, where $\alpha_{s}\neq 0$. Importantly, these biases carry substantial policy implications. Even with only 5% of isolated nodes, peer effect estimates are biased downward by over 35%. The bias seems not to be influenced by the share of isolated nodes.
Model 2 augments a single school fixed effect with a single global isolated-student dummy. Some researchers may adopt this specification to address the fact that isolated nodes have zero values for contextual variables. For instance, the share of female friends is zero for both isolated nodes and non-isolated nodes whose friends are exclusively male. To differentiate between these two cases, a dummy variable for isolated nodes can be incorporated into the model. While this approach corrects the bias observed in DGP B, it still results in a downward bias in DGP C. The bias is approximately 15% when the share of isolated nodes is 5%, increasing to 25% when 22% of nodes are isolated. This bias arises because of the assumption in DGP C that $\alpha_{s}$ varies across schools. [^13] As pointed out in Section 3.1, the bias depends on the sign of the covariance between the instrument and the omitted variable $\lambda\alpha_{s}\mathbf{g}_{s,i}\mathbf{1}_{n_{s}}$, as well as on the sign of the covariance between the instrument and the instrumented variable $\mathbf{G}_{s}\mathbf{y}_{s}$. It turns out that, by setting $\boldsymbol{\beta}=(-1,\penalty 10000\ -1.5)^{\prime}$ and $\boldsymbol{\gamma}=(-5,\penalty 10000\ 3)^{\prime}$—i.e., to the opposite of their actual values—so that the sign of the covariance between the instrument and $\mathbf{G}_{s}\mathbf{y}_{s}$ changes—Models 1 and 2 overestimate peer effects. We do not present these simulation results for brevity.
@@
Model 3, which is our preferred specification (Equation (9)), performs well across all DGPs. Yet, it is worth noting that the standard deviations of the estimates for DGPs A and B are higher in Model 3 compared to Model 2. This is because both models are suitable for these DGPs, whereas Model 3 is overparameterized when the true isolation gap is zero or constant.
@@
Model 1 Model 2 Model 3
Mean Sd Mean Sd Mean Sd
5% of Isolated Nodes
@@
$\lambda=0.7$ $0.700$ $0.015$ $0.700$ $0.015$ $0.701$ $0.022$
$\beta_{0,1}=1$ $1.000$ $0.047$ $1.000$ $0.047$ $1.000$ $0.048$
$\beta_{0,2}=1.5$ $1.500$ $0.084$ $1.000$ $0.047$ $1.500$ $0.084$
$\gamma_{0,1}=5$ $4.996$ $0.091$ $4.996$ $0.092$ $4.997$ $0.092$
$\gamma_{0,2}=-3$ 3.002-3.002 $0.114$ $4.996$ $0.092$ 3.005-3.005 $0.163$
@@
$\lambda=0.7$ $0.451$ $0.034$ $0.700$ $0.015$ $0.701$ $0.022$
$\beta_{0,1}=1$ $0.990$ $0.074$ $1.000$ $0.047$ $1.000$ $0.048$
$\beta_{0,2}=1.5$ $1.441$ $0.136$ $1.000$ $0.047$ $1.500$ $0.084$
$\gamma_{0,1}=5$ $4.541$ $0.186$ $4.996$ $0.092$ $4.997$ $0.092$
$\gamma_{0,2}=-3$ 5.756-5.756 $0.512$ $4.996$ $0.092$ 3.005-3.005 $0.163$